paula is 12 years older than his sister maria. four years from now bob will be two thirds as old as paula. how old are they now ?
Answer by stanbon(75887)

paula is 12 years older than his sister maria. four years from now bob will be two thirds as old as paula. how old are they now ?
------
Equations:
p = m + 12
b+4 = (2/3)(p+4)
----
You have 3 variables and only 2 equations.
Cheers,
Stan H.